A pancake recipe requires three and two thirds cups of milk to one cup of flour. If two and three quarters cups of milk is used, what quantity of flour will be needed, according to the recipe?

Answer by checkley79(3341)

(3 2/3)/1=(2 3/4)/x
(11/3)/1=(11/4)/x cross multiply.
11x/3=11/4 cross multiply.
4*11x=3*11
44x=33
x=33/44
x=3/4 cups of flour is needed.